OpenStudy (nicoleced):

help please. A=P+PRT solving for T i subtracted P from A and P itself so i now have T= A-PR. i feel like i skipped a step. but don't know how

OpenStudy (anonymous):

first solve for PRT...waht is it..?

OpenStudy (nicoleced):

PRT/PR?

OpenStudy (nicoleced):

since it's multiplication you do the opposite, right?

OpenStudy (anonymous):

no PRT = A-P

OpenStudy (anonymous):

then next.. divide both sides by PR to cancel it.. so what did u get..?

OpenStudy (anonymous):

what is ur T now..?

OpenStudy (nicoleced):

T = P/PR ?

OpenStudy (nicoleced):

wait no that's wrong!

OpenStudy (nicoleced):

T= a- p/pr ??

OpenStudy (nicoleced):

i don't know :(

OpenStudy (anonymous):

ur right...on that part...but u cn simplify it further..

OpenStudy (nicoleced):

Which answer was right?

OpenStudy (nicoleced):

@J-D

OpenStudy (anonymous):

T = \(\frac{ A-P }{ \Pr }\)

OpenStudy (nicoleced):

omg so you just flip it? cus thats the answer i had the first time but i thought it was wrong

OpenStudy (anonymous):

so is it clear now...?

OpenStudy (nicoleced):

yes thank you, i just need to practice more.